OpenStudy (anonymous):

i've been working on this for over a hour and i cannot figure out how to do this ): i will give a medal to best answer pls help! Create a quadratic polynomial function f(x) and a linear binomial in the form (x − a). Part 1. Show all work using long division to divide your polynomial by the binomial. Part 2. Show all work to evaluate f(a) using the function you created. Part 3. Use complete sentences to explain how the remainder theorem is used to determine whether your linear binomial is a factor of your polynomial function

OpenStudy (zzr0ck3r):

well you can construct this to make it easy on your self let our polynomial be (x-3)(x+4) = x^2+x-12 and your linear binomial to be (x-3) now we divide x^2+x-12 by x-3 we set this up like long division in grade school (x-3)|x^2+x-12 where | is the long division bar thingy so it would look like this

OpenStudy (zzr0ck3r):

you with me?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

now we divide the first term of (x-3) which is x by the first term of our polynomial, then we put that on top, then we multiply that by (x-3) and put that on bottom with a minus sign

OpenStudy (zzr0ck3r):

note x^2/x = x

OpenStudy (zzr0ck3r):

that should say "now we divide the first term of our polynomial by the first term of (x-3) which is x"...

OpenStudy (zzr0ck3r):

your not even here....

OpenStudy (zzr0ck3r):

gl

OpenStudy (anonymous):

im here.. it was glitching

OpenStudy (anonymous):

@zzr0ck3r